# RadioActive: 3D Radiological Interactive Segmentation Benchmark

### v2

@@ 4 Discussion and Conclusion
In this paper we introduced RadioActive  and with it, compared the performance of 2D and 3D interactive segmentation models in 3D medical imaging. We provide a holistic and transparent overview of the current state-of-the-art and highlight key findings that offer practical insights:
@@
    SAM 2 outperforms all medical 2D & 3D models: When prompting SAM2 through a realistic prompting scheme such as the Box Interpolation Sam 2 is superior to all medical 2D & 3D models. It even surpasses models after unrealistic iterative refinements.
@@
    Iterative Refinement is essential: The ability to iteratively refine segmentations significantly enhances model performance, particularly in challenging cases. Models that allow multiple rounds of corrections show better accuracy, making this feature crucial for clinical applications. For example, SegVol reached the highest performance in a static setting for point prompts, however, SAM-Med 3D Turbo is able to exceed SegVol given a few interactions, highlighting the importance of refinement.
@@
While the RadioActive benchmark addresses many issues, our work has limitations. Currently, we only approximate Human Effort. While the pragmatic approach provides valuable insights into model performance in relation to effort, it falls short of capturing the full complexity and practical challenges of real clinical applications. As future work, a comprehensive study involving clinicians is essential to assess the efficacy of different prompting strategies in real-world environments. Such a study should not only evaluate segmentation performance but also measure the time required for annotation, offering critical insights into the practical feasibility and efficiency of these models in clinical practice. Currently, the benchmark includes ten datasets, encompassing a variety of pathologies and selected anatomical structures in MRI and CT images. We acknowledge that the quantity of images in these datasets is relatively small—a common limitation for most rare lesion datasets—all large, however, most larger publicly available datasets have already been utilized in training the methods making this challenging. Another limitation of this work is that, given the rapid development of new interactive segmentation methods, very recent approaches, such as Wong et al. [41], are not yet included. However, we consider this a strength, as it highlights the need for a fair and open benchmark to keep up with ongoing advancements [26]. The published codebase is designed to seamlessly integrate new prompting strategies and models, enabling rapid adaptation and extension.
@@ Conclusion
The open-source nature of RadioActive  serves as a catalyst for interactive segmentation research in 3D medical imaging by enabling a fair and reproducible comparison between existing methods. Future work can build upon and extend this framework to investigate crucial questions in interactive segmentation. Through this, we believe RadioActive will contribute to improving real-world clinical application of interactive segmentation methods, facilitating a reduction of labor for medical professionals and accelerating meaningful clinical research.
@@
5 Model specifications
5.1 SAM
5.2 SAM2
5.3 MedSAM
5.4 SAM-Med2D
5.5 SAM-Med3D
5.6 SAM-Med3D Turbo
5.7 SegVol
@@ 6 Supported prompting schemes
All 2D and 3D bounding boxes were generated with perfect alignment around the target object. Since the impact of non-perfect bounding boxes has been extensively studied in previous works, we chose not to explore this additional degree of freedom [34, 27]. However, the open-source nature of our framework facilitates straightforward extensions to include a wider range of prompting scenarios.
7 Additional Results
@@
Difference between MRI und CT data
Previous prompts are essential for 2D model refinement
Previous prompts hinder 3D model refinement
For 3D models, Tab. 6 demonstrates that incorporating previous prompts does not enhance refinement performance beyond the first step. This observation aligns with the findings in Fig. 6 (right), where SAM-Med 3D and SAM-Med 3D Turbo show no improvement when provided with more than three prompts simultaneously.
Performance comparison between organs and pathological structures
Challenges of 3D patch-based prediction for larger structures
@@
1PPS SAM 1X 0.81 2.66 3.13 2.06 1.42 1.0 7.86 12.47 3.42 3.33 3.82
1PPS SAM2 1X 1.25 2.74 4.08 3.9 3.07 1.51 9.6 16.35 3.97 3.83 5.03
1PPS SamMed 2D 1X 10.72 26.3 24.55 28.83 22.76 9.71 30.83 31.24 26.05 20.31 23.13
2$\pm$±PPS SAM 2X 0.95 2.79 3.41 3.29 2.12 1.06 8.88 14.31 3.59 3.85 4.43
2$\pm$±PPS SAM2 2X 3.39 3.75 7.01 4.89 4.03 1.93 13.17 17.03 4.2 4.97 6.44
2$\pm$±PPS SamMed 2D 2X 11.88 28.09 27.62 33.29 24.07 11.25 32.54 31.5 28.27 21.27 24.98
2PPS SAM 2X 0.88 2.66 3.15 2.56 1.88 1.03 8.1 14.08 3.35 3.7 4.14
2PPS SAM2 2X 1.6 2.78 4.44 3.9 3.12 1.54 9.82 16.74 3.97 3.93 5.18
2PPS SamMed 2D 2X 11.63 29.9 28.9 32.88 24.38 11.34 37.84 36.41 32.51 22.39 26.82
3$\pm$±PPS SAM 3X 1.02 2.89 3.6 3.61 2.65 1.18 9.66 15.07 3.63 4.08 4.74
3$\pm$±PPS SAM2 3X 4.22 4.22 7.78 5.57 4.56 2.08 14.6 17.94 4.36 5.13 7.05
3$\pm$±PPS SamMed 2D 3X 12.83 31.8 31.72 36.4 25.85 12.62 40.42 36.91 34.12 23.29 28.6
3PPS SAM 3X 0.96 2.79 3.18 2.91 2.34 1.18 8.54 14.7 3.44 3.88 4.39
3PPS SAM2 3X 2.29 3.37 5.64 4.41 3.74 1.79 10.87 17.2 4.09 4.19 5.76
3PPS SamMed 2D 3X 11.95 31.34 31.8 33.61 24.9 12.15 43.26 40.16 35.62 23.07 28.79
5$\pm$±PPS SAM 5X 1.01 2.7 3.98 4.31 2.28 1.31 11.76 15.97 3.87 4.15 5.13
5$\pm$±PPS SAM2 5X 7.37 6.15 10.41 6.79 6.57 3.01 23.18 20.41 5.56 6.81 9.63
5$\pm$±PPS SamMed 2D 5X 13.95 34.18 35.63 39.39 27.19 14.01 46.49 41.05 37.83 24.5 31.42
5PPS SAM 5X 1.14 3.37 4.0 3.94 2.96 1.35 10.54 15.5 3.78 4.54 5.11
5PPS SAM2 5X 4.04 4.67 7.91 5.74 4.5 2.01 15.05 17.77 4.78 4.89 7.14
5PPS SamMed 2D 5X 12.26 32.17 34.55 33.08 25.24 12.66 50.88 45.0 38.34 23.22 30.74
10$\pm$±PPS SAM 10X 2.03 4.77 7.13 8.59 6.32 1.68 18.79 21.29 4.11 9.9 8.46
10$\pm$±PPS SAM2 10X 15.61 11.26 15.78 11.5 12.12 7.65 33.5 23.09 7.98 10.01 14.85
10$\pm$±PPS SamMed 2D 10X 15.02 35.62 39.67 42.49 28.27 15.95 52.98 46.17 40.97 25.26 34.24
10PPS SAM 10X 1.3 4.0 5.29 5.3 3.46 1.4 13.45 16.73 4.3 6.66 6.19
10PPS SAM2 10X 4.61 5.15 7.58 6.12 4.8 2.2 19.8 18.38 5.68 4.79 7.91
10PPS SamMed 2D 10X 12.03 31.1 36.39 30.38 24.22 13.24 58.07 51.41 38.98 22.1 31.79
Box PS SAM 2X 13.27 58.73 68.26 63.2 66.22 74.72 70.15 69.46 32.34 62.07 57.84
Box PS SAM2 2X 70.25 64.18 73.21 73.06 72.07 76.51 73.39 67.9 47.83 65.95 68.44
Box PS SamMed 2D 2X 28.2 46.08 57.04 64.45 46.44 41.91 62.07 63.59 45.62 41.51 49.69
Box PS MedSam 2X 40.63 47.39 55.5 60.4 45.73 46.43 67.75 70.23 48.82 46.13 52.9
@@
3P Inter SAM 3 0.84 3.6 3.49 2.22 1.67 1.03 8.28 12.67 3.48 4.56 4.18
3P Inter SAM2 3 1.38 3.67 4.75 4.18 3.41 1.63 10.32 17.01 4.17 5.18 5.57
3P Inter SamMed 2D 3 11.61 29.08 28.0 34.33 26.0 10.8 36.29 35.8 25.69 23.21 26.08
5P Inter SAM 5 0.84 3.6 3.5 2.17 1.7 1.02 8.3 12.61 3.64 4.5 4.19
5P Inter SAM2 5 1.38 3.68 4.73 4.1 3.37 1.63 10.44 16.98 4.4 5.13 5.58
5P Inter SamMed 2D 5 11.71 29.54 28.46 34.42 26.17 10.88 36.81 35.65 28.27 23.31 26.52
10P Inter SAM 10 0.84 3.61 3.48 2.21 1.71 1.03 8.26 12.64 3.66 4.49 4.19
10P Inter SAM2 10 1.38 3.67 4.76 4.14 3.43 1.62 10.39 17.01 4.5 5.1 5.6
10P Inter SamMed 2D 10 11.74 29.92 28.45 34.71 26.24 10.97 36.67 35.1 29.66 23.33 26.68
5P Prop SAM 7 1.09 3.49 3.6 1.78 1.7 0.99 7.84 13.7 3.39 4.32 4.19
5P Prop SAM2 7 3.77 3.75 4.59 3.47 3.13 1.6 8.96 17.52 3.78 4.79 5.54
5P Prop SamMed 2D 7 10.87 19.33 14.18 17.91 17.09 7.7 27.25 31.84 12.86 13.52 17.26
B Prop SAM 4 0.89 27.98 37.2 37.59 36.38 38.49 43.29 35.47 11.38 30.4 29.91
B Prop SAM2 4 3.82 33.77 43.31 40.82 41.08 40.88 46.54 37.39 19.37 31.68 33.87
B Prop SamMed 2D 4 2.82 23.78 30.17 36.12 25.15 22.56 37.98 45.6 19.1 21.0 26.43
B Prop MedSam 4 2.97 20.96 22.75 22.76 23.38 23.81 28.55 29.89 7.16 19.89 20.21
3B Inter SAM 6 13.13 54.04 63.34 58.64 64.26 71.35 65.52 61.78 25.58 59.23 53.69
3B Inter SAM2 6 69.76 57.59 67.71 69.08 69.06 73.11 68.35 63.24 37.21 63.26 63.84
3B Inter SamMed 2D 6 27.83 42.22 52.7 60.66 44.76 38.63 55.31 55.81 33.46 38.66 45.0
3B Inter MedSam 6 40.14 40.66 46.18 48.16 41.41 40.75 54.42 54.98 30.47 40.63 43.78
5B Inter SAM 10 13.25 56.3 66.81 62.58 65.76 73.63 69.5 68.25 31.33 61.54 56.9
5B Inter SAM2 10 70.13 61.16 71.53 72.57 71.32 75.52 72.72 67.25 46.01 65.42 67.36
5B Inter SamMed 2D 10 28.11 44.52 56.04 64.01 46.04 41.24 60.67 62.32 43.44 40.91 48.73
5B Inter MedSam 10 40.55 44.91 52.64 57.67 44.75 45.0 64.6 66.88 45.59 44.75 50.73
10B Inter SAM 20 13.27 57.71 67.81 63.17 66.12 74.35 70.02 69.21 32.24 61.98 57.59
10B Inter SAM2 20 70.25 63.0 72.65 73.05 71.94 76.12 73.19 67.62 47.48 65.83 68.11
10B Inter SamMed 2D 20 28.19 45.51 56.73 64.4 46.36 41.75 61.89 63.3 45.22 41.4 49.48
10B Inter MedSam 20 40.63 46.53 54.89 60.16 45.59 46.15 67.17 69.54 48.69 45.92 52.53
@@
1PPV SamMed 3D 1 1.92 10.95 21.19 29.12 13.28 16.66 56.06 48.55 42.81 23.63 26.42
1PPV SamMed 3D Turbo 1 5.03 18.34 30.08 18.38 10.75 34.93 28.51 19.26 39.26 17.06 22.16
1PPV SegVol 1 8.84 21.28* 31.49 25.49 2.45 32.66 61.77 52.46 29.09 25.85 29.14
1 center PPV SamMed 3D 1 2.07 12.15 24.11 26.9 15.11 19.64 72.66 53.24 50.4 26.51 30.28
1 center PPV SamMed 3D Turbo 1 5.18 27.34 46.07 34.46 15.91 46.38 82.98 59.49 63.83 26.98 40.86
1 center PPV SegVol 1 9.96 24.91* 38.49 31.36 3.17 33.92 71.01 50.67 28.21 30.73 32.24
2 center PPV SamMed 3D 2 1.87 11.51 23.15 24.45 13.14 18.19 71.11 45.59 40.95 25.36 27.53
2 center PPV SamMed 3D Turbo 2 5.33 26.62 45.71 33.26 15.75 46.77 84.88 70.5 69.56 27.49 42.59
2 center PPV SegVol 2 11.2 31.31* 47.51 58.45 11.57 52.36 75.08 52.66 33.4 32.45 40.6
3 center PPV SamMed 3D 3 1.77 11.29 22.45 23.41 12.19 17.31 70.31 46.24 36.77 24.85 26.66
3 center PPV SamMed 3D Turbo 3 5.16 26.48 43.81 30.93 15.28 46.56 85.75 68.52 69.67 27.19 41.94
3 center PPV SegVol 3 11.52 31.1* 50.08 57.51 18.76 53.46 73.04 58.22 45.52 33.95 43.32
5 center PPV SamMed 3D 5 1.73 10.95 21.86 21.29 11.21 16.04 66.47 43.95 34.78 22.73 25.1
5 center PPV SamMed 3D Turbo 5 5.09 25.75 42.74 26.85 14.67 46.5 85.92 64.7 70.11 26.6 40.89
5 center PPV SegVol 5 11.7 31.17* 49.08 52.47 25.4 52.85 61.6 45.66 44.16 33.48 40.76
10 center PPV SamMed 3D 10 1.72 10.54 20.87 20.78 10.4 14.54 61.01 41.78 27.3 20.16 22.91
10 center PPV SamMed 3D Turbo 10 5.06 24.9 40.04 21.51 13.1 46.03 86.05 61.5 68.76 25.45 39.24
10 center PPV SegVol 10 11.69 30.22* 45.24 47.32 26.43 51.68 42.04 26.69 41.88 32.96 35.61
3D Box SegVol 3 0.55 37.17* 68.11 69.72 63.21 50.13 89.95 79.98 72.13 49.45 58.04

### v3

@@ 4 Discussion and Conclusion
In this paper, we introduced RadioActive  and used it to compare the performance of 2D and 3D interactive segmentation models in 3D medical imaging. We provide a holistic and transparent overview of the current state-of-the-art and highlight key findings that offer practical insights, with additional findings detailed in the appendix Sec. 5:
@@
    SAM 2 outperforms all medical 2D & 3D models: When prompting SAM2 through a realistic prompting scheme such as the Box Interpolation SAM2 is superior to all medical 2D & 3D models. Future studies should benchmark against SAM2 with 3-box interpolation and scribble refinement as the new gold standard.
@@
    Iterative Refinement is essential: The ability to iteratively refine segmentations significantly enhances performance, particularly in challenging cases. Models that allow multiple rounds of corrections show better accuracy, making this feature crucial for clinical applications. For example, SegVol reached the highest performance in a static setting for point prompts, however, SAM-Med 3D Turbo can match SegVol given a few interactions.
@@
While the RadioActive benchmark addresses many issues, our work has limitations. Currently, we only approximate Human Effort, providing valuable insights but not fully capturing the complexity of real clinical applications. A crucial next step is a comprehensive study with clinicians to assess the efficacy of different prompting strategies in real-world settings. Beyond segmentation performance, such a study should measure annotation time to evaluate practical feasibility. Ultimately, integrating the best prompting methods into an actual annotation toolkit will be key to improving efficiency in clinical practice. While the datasets used in our study contain relatively few images, our benchmark includes more datasets than any previous study (ten datasets) and covers a wide range of pathologies and anatomical structures in MRI and CT. Moreover, we denote that most larger publicly available datasets have already been used for training, leaving limited options to choose datasets from. Another limitation is that, given the rapid development of new interactive segmentation methods, very recent approaches are not yet included. However, this also highlights the need for a fair and open benchmark to keep up with ongoing advancements [29]. The published codebase is designed to seamlessly integrate new prompting strategies and models, enabling rapid adaptation and extension.
@@ Conclusion
RadioActive  serves as a catalyst for interactive segmentation research in 3D medical imaging by enabling a fair and reproducible comparison between existing methods by its open-source nature. Future work can build upon and extend this framework to investigate crucial questions in interactive segmentation. Through this, RadioActive will contribute to improving real-world clinical application of interactive segmentation methods, facilitating a reduction of labor for medical professionals and accelerating clinical research.
@@
5 Additional Findings
* Refer to caption
Difference between MRI und CT data
Previous prompts are essential for 2D model refinement.
SamMed 3D and SamMed 3D Turbo react less to refinement prompts.
SegVol does not support iterative refinement since it does not accept mask prompts. SamMed 3D and SamMed 3D Turbo are trained exclusively with a single point prompt as input, which we believe makes iterative refinement more challenging. As with the issue mentioned above for 2D models, this limitation can lead to losing track of the target structure.
Performance comparison between organs and pathological structures
Challenges of 3D patch-based prediction for larger structures
ScribblePromt performs better with points than with boxes.
We believe this is because, despite being trained on various prompt types, the method is primarily designed for scribble-based interactions, as its name suggests. Since points are more similar to scribbles, they achieve better performance than boxes. We did not include optimized scribble prompts as the initial interaction since other methods do not support them. Notably, ScribblePrompt reacts strongly to multiple points, which may resemble a scribble in interaction style, leading to a significant performance improvement.
SAM reacts less to additional promts than SAM2.
Both for multiple initial point promts, as shown in Tab. 2, and for iterative refinement, as shown in Tab. 5, SAM2 improves more with additional prompts.
6 Current State of Independent Benchmarkings
SAM’s introduction spurred numerous efforts to use it for medical applications, leading to a wave of studies evaluating its performance on medical data. Early benchmarks primarily compared SAM with non-interactive models [14], [38] or outdated interactive methods [30]. More recently, an influx of medical-specific SAM adaptations has emerged [27, 4, 43, 11, 37, 45]. While these studies present comparisons with baseline methods, they tend to focus on evaluation settings where their approach excels. Independent benchmarking remains essential for obtaining an unbiased and comprehensive assessment of relative performance [29], particularly given the opaque and insufficient evaluation practices observed in some of these works (P3).
Currently, most existing benchmarkings focus on analyzing the behavior of only one or two models, mostly SAM and SAM2 [48, 1, 19, 38], or only evaluating 2D images [5, 1, 19, 38]. While the comprehensive 2D medical benchmarking study by Chen et al. [5] incorporates multiple recent methods, it has several limitations. Notably, it evaluates only three external datasets, two of which (SegThor and ISLES2022) were used in the training of both MedSAM and SAM-Med2D. Additionally, the evaluation focuses on larger, homogeneous target structures, lacking smaller, more challenging pathological structures that are difficult to delineate.
To our knowledge, no existing work has specifically addressed a realistic scenario for 3D interactive segmentation. This highlights a gap in the current benchmarking literature, which has yet to catch up with recent advancements in interactive medical segmentation. Given the rapid development of new 2D and 3D models expected in the coming months and years, this remains a significant challenge. Our independent benchmarking not only updates the field but also introduces an open-source framework that enables future benchmarking efforts to be conducted more efficiently, transparently, and comparably.
@@ 7 Supported prompting schemes
All 2D and 3D bounding boxes were generated with perfect alignment around the target object. Since the impact of non-perfect bounding boxes has been extensively studied in previous works [38, 30], we only explore this additional degree of freedom for the proposed new gold standard Sam2 with 3 Box interpolation in Tab. 2.
@@
Figure 9: Realistic Scribble Refinement. When the previous segmentation overestimates the target, we identify the non-axial slice with the highest number of false positives and generate a contour around the ground truth object at a 2-pixel distance. A 60% segment of this contour is then selected, using all false positive points along it as negative refinement prompts for 2D models, while a random FP point of the curve is chosen for 3D models. Conversely, when the segmentation underestimates the target, we perform a connected component analysis on false negative voxels and generate a ”scribble” through the center of the largest false negative component. Each axial point along this scribble serves as a new positive prompt for 2D models and one random point of this scribble for 3D models.
8 Model specifications
8.1 SAM
8.2 SAM2
8.3 MedSAM
8.4 SAM-Med2D
8.5 SAM-Med3D
8.6 SAM-Med3D Turbo
8.7 SegVol
8.8 ScribblePromt
Each image slice is normalized between 0 and 1 and reshaped to a fixed 128x128 image size. Inference was restricted to slices containing foreground. The predicted slices were inverse transformed to the original coordinate system, and metrics were computed in the original image space. Notably, we apply ScribblePrompt only on axial slices, but it was originally trained on all views.
@@ 8.8 ScribblePromt
SAM 1PPS 1X NaN 3.42 3.06 4.21 1.43 0.92 NaN 15.4 28.52 4.89 7.73
SAM2 1PPS 1X 3.36 4.99 4.41 5.91 3.77 1.62 15.64 22.36 45.86 6.79 11.47
SamMed 2D 1PPS 1X 18.28 33.98 44.01 41.44 35.97 18.79 59.09 49.33 62.66 24.28 38.78
ScribblePrompt 1PPS 1X 25.29 32.9 47.38 49.09 26.43 4.37 49.73 45.4 55.47 35.12 37.12
SAM 2PPS 2X 1.98 3.59 3.16 6.22 2.18 0.99 12.29 17.34 28.95 5.43 8.21
SAM2 2PPS 2X 4.4 5.83 5.8 6.67 4.41 1.62 17.82 22.35 48.72 7.63 12.53
SamMed 2D 2PPS 2X 17.99 37.44 44.99 44.82 39.0 20.38 68.51 54.9 67.2 27.13 42.24
ScribblePrompt 2PPS 2X 37.15 43.31 60.66 67.58 39.55 14.9 59.35 51.42 65.31 43.83 48.31
SAM 2$\pm$±PPS 2X 3.47 3.91 3.8 7.47 2.19 1.31 15.5 19.88 38.33 6.93 10.28
SAM2 2$\pm$±PPS 2X 11.24 10.86 12.73 11.0 6.42 2.87 38.04 29.13 54.47 11.18 18.79
SamMed 2D 2$\pm$±PPS 2X 19.62 36.56 47.23 45.41 37.39 21.44 61.12 50.44 64.5 25.06 40.88
ScribblePrompt 2$\pm$±PPS 2X 34.08 37.68 50.98 64.55 35.92 24.43 54.23 47.68 57.72 35.07 44.23
SAM 3PPS 3X 2.07 4.0 3.38 7.22 2.92 1.19 14.27 19.36 31.1 5.84 9.13
SAM2 3PPS 3X 6.19 8.26 9.35 9.29 5.72 1.99 28.66 24.23 53.18 8.32 15.52
SamMed 2D 3PPS 3X 17.61 38.93 45.34 44.74 38.86 21.14 73.81 58.24 69.45 27.9 43.6
ScribblePrompt 3PPS 3X 43.23 50.1 67.63 74.49 46.89 32.53 68.6 59.62 70.65 48.07 56.18
SAM 3$\pm$±PPS 3X 3.82 4.33 4.12 9.38 2.87 1.42 17.77 21.84 37.91 7.13 11.06
SAM2 3$\pm$±PPS 3X 13.75 12.49 15.33 14.66 8.0 3.24 44.98 30.34 57.62 11.75 21.22
SamMed 2D 3$\pm$±PPS 3X 19.59 39.52 47.98 47.76 40.27 22.82 71.01 56.28 68.38 27.77 44.14
ScribblePrompt 3$\pm$±PPS 3X 43.19 47.89 62.72 73.52 48.57 43.65 65.73 58.14 66.71 44.67 55.48
SAM 5PPS 5X 4.19 5.56 5.94 11.05 4.12 1.52 22.86 23.88 40.75 7.88 12.77
SAM2 5PPS 5X 8.68 12.59 15.58 14.89 7.32 2.68 48.72 27.08 55.08 9.76 20.24
SamMed 2D 5PPS 5X 17.06 39.74 45.49 42.65 37.08 21.57 78.73 62.38 71.48 27.74 44.39
ScribblePrompt 5PPS 5X 48.0 57.41 72.75 77.77 54.06 51.98 76.77 69.31 75.79 52.39 63.62
SAM 5$\pm$±PPS 5X 5.17 4.43 5.74 12.05 2.87 1.78 24.41 27.98 42.64 8.51 13.56
SAM2 5$\pm$±PPS 5X 20.17 19.59 24.07 25.43 12.91 6.17 62.15 38.99 64.65 17.33 29.15
SamMed 2D 5$\pm$±PPS 5X 20.05 41.94 49.86 49.43 40.72 24.75 76.75 60.72 70.94 28.65 46.38
ScribblePrompt 5$\pm$±PPS 5X 50.01 56.37 68.66 77.95 58.49 58.6 69.59 64.63 72.25 49.98 62.65
SAM 10PPS 10X 5.5 6.64 8.14 15.7 4.88 2.14 31.65 28.55 47.06 16.5 16.68
SAM2 10PPS 10X 6.67 13.62 14.5 16.1 8.13 3.45 55.1 31.33 54.52 9.38 21.28
SamMed 2D 10PPS 10X 15.94 38.7 44.96 38.15 32.23 22.22 82.57 67.07 71.92 25.7 43.95
ScribblePrompt 10PPS 10X 51.46 65.06 76.14 79.34 60.15 60.46 77.17 73.45 80.27 56.28 67.98
SAM 10$\pm$±PPS 10X 9.75 8.0 10.6 23.07 9.61 2.02 28.43 35.11 45.62 22.77 19.5
SAM2 10$\pm$±PPS 10X 28.7 29.26 36.21 40.99 24.98 18.53 71.71 46.53 72.77 24.55 39.42
SamMed 2D 10$\pm$±PPS 10X 21.05 43.97 51.95 50.31 38.98 27.87 81.8 66.4 72.97 28.31 48.36
ScribblePrompt 10$\pm$±PPS 10X 55.75 66.81 76.11 82.16 68.4 69.71 72.94 68.89 78.28 57.27 69.63
SAM Box PS 2X 75.93 65.2 82.79 83.68 76.67 80.97 89.06 84.17 90.23 71.01 79.97
SAM2 Box PS 2X 78.0 72.01 86.57 88.5 81.58 85.21 91.15 89.01 89.0 75.28 83.63
SamMed 2D Box PS 2X 40.41 56.79 79.73 72.84 58.12 59.11 88.97 82.2 82.71 48.01 66.89
ScribblePrompt Box PS 2X 14.68 29.2 53.97 65.9 24.61 12.4 25.53 23.03 65.46 10.39 32.52
MedSam Box PS 2X 58.22 69.98 72.99 76.98 69.9 68.18 92.87 89.74 68.1 55.95 72.29
@@
SAM 3P Inter 3 2.06 3.65 3.06 4.81 1.41 0.91 11.05 15.45 14.81 6.63 6.38
SAM2 3P Inter 3 3.76 5.39 4.5 6.52 3.74 1.55 15.6 22.57 23.96 9.22 9.68
SamMed 2D 3P Inter 3 19.13 34.07 44.69 47.93 42.86 20.63 NaN 49.54 55.82 27.38 38.01
ScribblePrompt 3P Inter 3 27.69 34.7 56.69 59.87 28.6 4.58 50.46 45.5 52.22 41.4 40.17
SAM 5P Inter 5 2.05 3.57 3.07 4.84 1.41 0.91 11.04 15.43 20.65 6.7 6.97
SAM2 5P Inter 5 3.77 5.37 4.52 6.83 3.76 1.55 15.61 22.59 32.41 9.12 10.55
SamMed 2D 5P Inter 5 19.27 34.86 44.91 48.21 43.35 20.77 64.46 49.73 60.7 27.6 41.39
ScribblePrompt 5P Inter 5 27.93 35.68 57.27 59.45 28.82 4.58 50.77 45.39 57.67 41.49 40.9
SAM 10P Inter 10 2.05 3.6 3.06 4.8 1.41 0.92 11.05 15.46 25.25 6.48 7.41
SAM2 10P Inter 10 3.77 5.38 4.51 6.72 3.74 1.55 15.71 22.5 40.64 8.7 11.32
SamMed 2D 10P Inter 10 19.33 35.48 44.75 48.18 43.56 20.73 64.51 48.64 62.7 27.67 41.56
ScribblePrompt 10P Inter 10 28.01 36.45 57.6 59.56 28.84 4.61 50.95 44.81 59.56 41.52 41.19
SAM 5P Prop 7 3.91 3.93 3.2 4.17 1.82 0.91 11.55 14.37 12.74 6.31 6.29
SAM2 5P Prop 7 7.82 6.37 4.66 5.57 4.44 1.63 14.89 19.61 17.25 7.93 9.02
SamMed 2D 5P Prop 7 15.5 26.04 30.54 30.25 28.69 17.98 58.1 47.94 42.37 16.95 31.44
ScribblePrompt 5P Prop 7 41.08 31.39 42.63 38.61 24.88 4.85 43.78 33.65 32.1 35.27 32.82
SAM B Prop 4 4.77 37.71 53.06 65.66 46.49 43.74 69.39 50.39 76.49 40.41 48.81
SAM2 B Prop 4 4.76 42.35 60.54 67.65 50.53 49.55 76.06 67.11 74.4 39.23 53.22
SamMed 2D B Prop 4 5.0 35.97 60.33 52.7 35.09 37.46 74.3 69.5 65.06 26.43 46.19
ScribblePrompt B Prop 4 3.43 12.27 21.49 46.45 12.08 5.16 7.69 3.24 13.99 3.79 12.96
MedSam B Prop 4 4.8 39.38 38.71 45.48 45.54 42.95 57.25 55.32 15.47 29.11 37.4
SAM from 3D Box 3 69.67 47.56 60.62 67.56 59.08 63.77 76.43 57.28 50.0 54.28 60.63
SAM2 from 3D Box 3 71.05 51.5 62.5 66.23 62.89 66.93 76.65 60.62 41.27 56.31 61.6
SamMed 2D from 3D Box 3 37.7 46.19 65.53 58.18 48.06 53.71 75.76 64.52 47.57 39.05 53.63
ScribblePrompt from 3D Box 3 15.45 30.41 56.52 65.43 33.7 15.38 36.47 34.42 56.19 12.61 35.66
MedSam from 3D Box 3 54.94 59.43 68.82 71.48 68.14 67.02 85.53 73.18 31.32 48.94 62.88
SAM 3B Inter 6 75.41 60.53 80.06 81.65 75.75 77.85 86.54 76.3 62.48 67.48 74.41
SAM2 3B Inter 6 77.29 65.06 80.2 82.63 78.6 80.43 86.29 78.43 58.8 71.41 75.91
SamMed 2D 3B Inter 6 39.98 52.3 75.68 68.33 56.34 56.67 80.76 69.65 58.4 44.42 60.25
ScribblePrompt 3B Inter 6 14.16 24.84 44.69 54.56 18.72 9.78 17.79 9.25 32.14 7.09 23.3
MedSam 3B Inter 6 57.46 60.45 61.44 65.3 64.57 61.43 72.82 67.6 38.8 49.05 59.89
SAM 5B Inter 10 75.8 63.21 82.1 83.59 76.32 80.07 89.29 83.54 86.57 70.25 79.08
SAM2 5B Inter 10 77.83 69.05 84.69 87.5 80.77 83.83 90.81 87.13 82.67 74.55 81.88
SamMed 2D 5B Inter 10 40.34 55.13 78.8 72.31 57.63 58.42 87.67 80.17 79.12 47.26 65.68
ScribblePrompt 5B Inter 10 14.58 27.84 51.5 63.74 23.54 11.88 23.3 16.74 54.92 9.48 29.75
MedSam 5B Inter 10 58.09 66.79 69.7 74.97 68.81 66.75 89.11 84.64 59.6 54.28 69.27
SAM 10B Inter 20 75.93 64.49 82.38 83.54 76.57 80.61 89.08 84.21 89.99 70.84 79.76
SAM2 10B Inter 20 77.99 70.88 85.95 88.26 81.44 84.8 91.02 88.61 88.28 75.08 83.23
SamMed 2D 10B Inter 20 40.41 56.3 79.4 72.63 58.04 58.99 88.79 81.88 82.37 47.84 66.67
ScribblePrompt 10B Inter 20 14.68 28.75 53.45 65.72 24.56 12.33 25.12 21.38 63.31 10.26 31.96
MedSam 10B Inter 20 58.22 68.9 72.2 76.64 69.77 67.85 92.23 88.6 66.52 55.74 71.67
@@
SegVol 1 center PPV 1 9.96 24.91 38.49 31.36 3.17 33.92 71.01 50.67 28.21 30.73 32.24
SamMed 3D Turbo 1 center PPV 1 5.18 27.34 46.07 34.33 15.91 46.38 82.95 59.5 63.75 26.98 40.84
SamMed 3D 1 center PPV 1 2.07 12.15 24.06 27.16 15.11 19.64 72.66 53.24 50.38 26.51 30.3
SegVol 1PPV 1 8.84 21.28 31.49 25.49 2.45 32.66 61.77 52.46 29.09 25.85 29.14
SamMed 3D Turbo 1PPV 1 5.03 18.34 30.08 18.38 10.75 34.93 28.51 19.26 39.26 17.06 22.16
SamMed 3D 1PPV 1 1.92 10.95 21.19 29.12 13.28 16.66 56.06 48.55 42.81 23.63 26.42
SegVol 2 center PPV 2 11.2 31.31 47.51 58.45 11.57 52.36 75.08 52.66 33.4 32.45 40.6
SegVol 2PPV 2 22.53 32.08 44.9 51.82 9.01 53.99 70.19 57.02 36.2 33.26 41.1
SegVol 3 center PPV 3 11.52 31.1 50.08 57.51 18.76 53.46 73.04 58.22 45.52 33.95 43.32
SegVol 3PPV 3 25.46 37.33 51.77 62.41 16.79 59.36 77.44 62.83 47.21 36.57 47.72
SegVol 5 center PPV 5 11.7 31.17 49.08 52.47 25.4 52.85 61.6 45.66 44.16 33.48 40.76
SegVol 5PPV 5 26.07 41.98 61.5 65.13 27.73 62.78 82.23 71.76 59.96 40.47 53.96
SegVol 10 center PPV 10 11.69 30.22 45.24 47.32 26.43 51.68 42.04 26.69 41.88 32.96 35.61
SegVol 10PPV 10 26.09 46.51 68.3 66.98 38.92 64.67 84.18 76.73 70.49 44.74 58.76
SegVol 3D Box 3 0.55 37.17 68.11 69.72 63.21 50.13 89.95 79.98 72.13 49.45 58.04